OpenStudy (anonymous):

Need help with the following questions:

OpenStudy (anonymous):

1) If f(x)=x^4-14x^3+72X^2+9x-10 find the largest interval on which f(x) is concave downward. If we write the interval as (a,b), then a=___ and b=___ 2) If g(x)=x^4-12x^3+30x^2+9x-3 find the largest interval on which g(x) is concave upward. If we write the interval as (-infinity, a) U (b, infinity), then a=___ and b=___ 3) If f(x)=xe^6x find the largest interval on which f(x) is concave upward. If we write the interval as (a, infinity), then a=___

OpenStudy (anonymous):

do i take the derivative of each twice to get these answers?

OpenStudy (anonymous):

@satellite73 @amistre64 @KingGeorge @.Sam. @Mertsj @saifoo.khan

OpenStudy (anonymous):

someone help? i dont understand this stuff completely yet

OpenStudy (kinggeorge):

You would take the second derivative twice, and find the zeros to determine where it switches concavity.

OpenStudy (kinggeorge):

Then plug in values in the different intervals you get to find if the concavity is positive or negative in that interval.

OpenStudy (anonymous):

show me the work for one? i get the gist of it, just wanna make sure

OpenStudy (kinggeorge):

Let's take the first one then. You have\[f(x)=x^4-14x^3+72x^2+9x-10\]so\[f'(x)=4x^3-42x^2+144x+9\]\[f''(x)=12x^2-84x+144=12(x^2-7x+12)\]Now you set that equal to zero, and you get\[0=12(x^2-7x+12)=(x-3)(x-4)\]So you have zeros at \(x=3\), \(x=4\).

OpenStudy (kinggeorge):

Finally, plug in some number for \(x\) that is less than three, one that is between 3 and 4, and one that is greater than 4, and check which one(s) are negative. In this case, only the one between 3 and 4 is negative, so that must be greatest interval on which \(f(x)\) is concave downward.

OpenStudy (anonymous):

what happened to the 12 up there. did you divide by 12 on both sides?

OpenStudy (kinggeorge):

Yup.

OpenStudy (anonymous):

so the answer would be like 3.5

OpenStudy (kinggeorge):

You want an interval, not a single number, so your interval should be \((3, 4)\). Be careful that it doesn't include \(3\) or \(4\), since the concavity is zero there.

OpenStudy (anonymous):

o yeah duh. i wasnt thinking straight.

OpenStudy (anonymous):

when we plug the numbers less, in between, and greater than, we plug them in the original function, correct?

OpenStudy (kinggeorge):

Not if you want to find the concavity. You would want to plug them into the second derivative.

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

i know prob 3 is similar but could u help me with that as well...

OpenStudy (anonymous):

3) If f(x)=xe^6x find the largest interval on which f(x) is concave upward. If we write the interval as (a, infinity), then a=___

OpenStudy (kinggeorge):

First we need to find the second derivative. Using the product rule and chain rule\[f(x)=xe^{6x}\]\[f'(x)=6xe^{6x} + e^{6x}=e^{6x}(6x+1)\]\[f''(x)=6e^{6x}+36xe^{6x}\]Now you find the zeros.

OpenStudy (kinggeorge):

So\[0=6e^{6x}+36xe^{6x}=e^{6x}+6xe^{6x}=e^{6x}(1+6x)\] since \(e^{6x}\) is never 0, the only zero is at \(1+6x=0 \to x=-1/6\)

OpenStudy (anonymous):

so whats the final answer im confused

OpenStudy (anonymous):

calc is so complex :(

OpenStudy (kinggeorge):

That's where the concavity switches. So if we plug in \(x=0\), we have that the concavity is upward from \((-1/6, \infty)\)

OpenStudy (anonymous):

i put that as my answer and it said it was wrong

OpenStudy (anonymous):

u sure u did it right?

OpenStudy (anonymous):

im confused by some of ur work above ^^^

OpenStudy (kinggeorge):

You have to notice that the question is calling for a single value \(a\), not an interval. Since it gives you the interval as \((a, \infty)\), your answer should be \(a=-1/6\).

OpenStudy (anonymous):

i know. it wont accept that

OpenStudy (kinggeorge):

I think I did the second derivative wrong. It should actually be\[f''(x)=12e^{6x}(3x+1)\]

OpenStudy (kinggeorge):

So \(a=-1/3\) instead.

OpenStudy (anonymous):

ok thats its. can u rewrite it all correctly?

OpenStudy (anonymous):

*it

OpenStudy (kinggeorge):

Give me a minute.

OpenStudy (anonymous):

sure

OpenStudy (kinggeorge):

\[f(x)=xe^{6x}\]\[f'(x)=6 xe^{6x} +e^{6x} = e^{6x}(6x+1)\]\[f''(x)=6 e^{6x} (6 x+1)+6 e^{6x}=12e^{6x}(3x+1)\]Since \(12e^{6x}\) is never zero, we know that \((3x+1)=0\) So \(x=-1/3\).

OpenStudy (anonymous):

THANKS

OpenStudy (kinggeorge):

You're welcome. Sorry it took so long.

OpenStudy (anonymous):

its fine. been doing other problems while waiting.